OpenStudy (anonymous):

Factor the trinomial completely. 20x^2-7x-6

OpenStudy (anonymous):

(5x+2)(4x-3)

OpenStudy (anonymous):

I can't really explain how. It was guess and check.

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

I see that this is second same question :) I'll help you make it clear, so you can easily do any of them by yourself :) You can ALWAYS use formula! http://en.wikipedia.org/wiki/Quadratic_equation For x1 you will get 3/4, and for x2 you will get -2/5 So you have: 1) x=3/4 ALL you do is -> x-3/4=0 2) x=-2/5 ALL you do is -> x+2/5=0 So result you want is (x-3/4)*(x+2/5) Clear enough? :) Hope I helped! Good luck

OpenStudy (anonymous):

or (20x+8)(x-15/20)

OpenStudy (anonymous):

I started with 2 products of the first term (5 and 4). Then I chose 2 products of the last term (2 and 3) I set it up like this (5x 3)(4x 2) I knew i needed the sum to be -7 so the only way that would work is 15-8. So I flipped it to (5x 2)(4x 3) then I filled in the signs

OpenStudy (anonymous):

Thank you @nickersia

OpenStudy (anonymous):

thank you @masterxnonos

OpenStudy (anonymous):

I like the quadratic formula tho. Their answer matched mine if you just get rid of the fractions. I'd stick to that. easy